40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 5 20:37:34 2011, jamie, Summary, General, Summary of today's in-vacuum work 
    Reply  Sat Aug 6 03:10:05 2011, Suresh, Summary, General, Summary of today's in-vacuum work ETMX_OSEMs.pngAfter_osem_adjustment.JPGBefore_osem_adj.JPG
       Reply  Sat Aug 6 13:38:02 2011, rana, Summary, General, Summary of today's in-vacuum work 
          Reply  Sun Aug 7 11:10:54 2011, Suresh, Summary, General, OSEM positioning 
    Reply  Sun Aug 7 14:01:58 2011, kiwamu, Update, Green Locking, X green beam re-aligned 
       Reply  Sun Aug 7 18:41:12 2011, Jenne, Update, Green Locking, X green beam re-aligned 
Message ID: 5135     Entry time: Sun Aug 7 18:41:12 2011     In reply to: 5133
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: X green beam re-aligned 

Quote:

The ETMY suspension tower is in a good place.

Quote from #5127

Kiwamu will work on the green alignment over the weekend.  Assuming everything works out, we'll try the same procedure on ETMY on Monday.

 

 You meant ETMX, right?  ETMY still hasn't been touched.

KI : sorry, I meant ETMX. I fixed the entry.

ELOG V3.1.3-